OpenStudy (anonymous):

14 piece of cable is cut into 2 pieces. one piece is 4 m longer than the other. how long are the pieces?

OpenStudy (kropot72):

Let the length of the shorter piece of cable = x meters The length of the longer piece is x + 4 meters x + x + 4 = 14 2x + 4 = 14 2x = 14 - 4 = 10 x = 10/2 meters and the length of the longer piece is 10/2 + 4 meters.